Please help me solve this using substitution and elimination { { { -8x+9y=69 3y-6x=33 } } }
Answer by addingup(3677)

We have two equations:
-8x+9y = 69
and
3y-6x = 33
Rearrange the terms on this last equation:
-6x+3y = 33 Multiply both sides of this equation times 3:
-18x+9y = 99
----------------------------------------
Now let's subtract this last equation from the first:
-8x+9y = 69
-
-18x+9y = 99 (Note: We have a - -: -8x - -18x = -8x + 18x)
--------------
10x = -30
x= -3
Now let's find the value of y:
-8(-3)+9y= 69
24+9y = 69
9y = 45
y = 5
Finally, we'll check to see if we have the right answer. The right answer must make the two equations true (where one side equals the other, that's why it's called an equation)
-8x+9y = 69; -8(-3)+9(5)= 69; 24+45 = 69; 69 = 69
and
3y-6x = 33 3(5)-6(-3)= 33; 15+18 = 33 Both equations are true, we have the correct answer.